OpenStudy (anonymous):

find the solutions of the system. 10. y = x² + 3x – 4 y = 2x + 2 (–3, 6) and (2, –4) (–3, –4) and (2, 6) (–3, –4) and (–2, –2) no solution 11. y = x² – 2x – 2 y = 4x + 5 (–1, 1) and (–7, –23) (–1, 1) and (7, 33) (–1, 33) and (7, 1) no solution

OpenStudy (anonymous):

you have: \[ \LARGE y = x^2 + 3x – 4\] \[ \LARGE y =2x+2 \] substitute the second equation into the first one... \[ \LARGE 2x+2 = x^2 + 3x – 4\] \[ \LARGE 2x-3x-x^2+4+2 = 0\] \[ \LARGE-x^2-x+6=0\] multiply both sides by -1 \[ \LARGE x^2+x-6=0\] now you have a quadratic equation there... use its formula: \[\LARGE {x_{1/2}} = \frac{{ - b \pm \sqrt {{b^2} - 4ac} }}{{2a}}\] substitute and there you go :)

OpenStudy (anonymous):

so what the answer

OpenStudy (anonymous):

0

OpenStudy (anonymous):

so no solution

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

(–3, 6) and (2, –4) (–3, –4) and (2, 6) (–3, –4) and (–2, –2) no solution

OpenStudy (anonymous):

hey wait... there are Two solutions !!! \[\LARGE x_1/2=\frac{-1\pm\sqrt{1^2-4\cdot (-6)}}{2 }\] \[\LARGE x_1/2=\frac{-1\pm\sqrt{1+24 }}{2 }\] \[\LARGE x_1/2=\frac{-1\pm5}{2 }\] !!!

OpenStudy (anonymous):

hold on... I think I just Got dizzy ! O_O

OpenStudy (anonymous):

what a grind lol

OpenStudy (anonymous):

(-3,-4) and (2,6) this is the solution...

OpenStudy (anonymous):

I haven't got dizzy yet :F

OpenStudy (anonymous):

lol^

OpenStudy (anonymous):

^^